COMBINED EXPERIMENTAL AND THEORETICAL STUDIES ON THE VIBRATIONAL SPECTRA OF SOME QUINOLINECARBOXALDEHYDE MOLECULES-I

\underline{MUSTAFA KUMRU}, Department of Physics, Faculty of Arts and Sciences, Fatih University, 34500 Büyükçekmece, Istanbul, Turkey, mkumru@fatih.edu.tr.

Combined experimental and theoretical studies have been performed on the structure and vibrational spectra (FT-IR, FT-Raman, Dispersive Raman spectra including far region) of some quinolinecarboxaldehyde molecules. Hartree-Fock (HF) and density functional B3LYP calculations have been employed with the 6-311++G(d,p) basis set for investigating the structural and spectroscopic properties of the cis and trans conformers of quinolinecarboxaldehyde. The B3LYP frequencies are closer to the experimental frequencies than the HF frequencies, but scaled frequencies of both HF and B3LYP agree almost perfectly with the experimental frequencies. The cis conformer has been found more stable than the trans conformer. The scaled vibrational frequencies of cis conformer also agree slightly better than those of the trans conformer with the experimental frequencies.

1. V. Kçük, A. Altun, M. Kumru, Spectrochimica Acta Part A 85(2012)92-98

2. M. Kumru, V. Küçük, T. Bardakç, Spectrochimica Acta Part A 90(2012)28-34 We thank the Turkish Scientific and Technical Research Council (TÜBÍTAK) for their financial support through National Postdoctoral Research Scholarship Programme and Scientific Research Fund of Fatih University under the project number P50011001 G (1457).